OpenStudy (anonymous):

Help Please....... Find the derivative of the function by using the definition.

OpenStudy (anonymous):

\[y = \frac{ 1 }{ 4 }x^4 + \frac{ 1 }{ 3 }x^3 - 1\]

OpenStudy (anonymous):

Deriv of x^n n(x^(n-1)) In your case n =4

OpenStudy (anonymous):

4(1/4)(x^(4-1) + 3(1/3)(x^(3-1)) + 0 derivative =x^3 + x^2 Derivative of a constant is zero